Suppose that the consumer’s utility function is given by 1/ 2 1/ 4 U  10X Y . If price of good X and Y are 4 and 2

respectively and income constraint is Birr 100.

a. The demand equation for X and Y.

b. The utility maximizing levels of X and Y.

c. The maximum utility.

d. Compute both X ,Y Y ,X MRS and MRS . Is there any difference between the two?


1
Expert's answer
2022-01-23T15:55:26-0500

"U(10X^\frac {1}{2}Y^\frac{1}{4})"

"P_x=4"

"P_y=2"

Income= 100

a)

"\frac{Mu_x}{Mu_y}=\frac{P_x}{P_y}"


"Mu_x=5x^\frac{-1}{2}Y^\frac{1}{4}"


"Mu_y=2.5x^\frac{1}{2}Y^\frac{-3}{4}"


"\frac{5x^\frac{-1}{2}Y^\frac{1}{4}}{2.5x^\frac{1}{2}Y^\frac{-3}{4}}=\frac{P_y}{P_x}"


"\frac{5Y}{2.5X}=\frac{P_x}{P_y}"


"Y=\frac{2.5XP_x}{5P_y}"


="\frac{0.5XP_x}{P_y}" ...........(i)

"X=\frac{5P_yY}{2.5P_x}"


"X=\frac{2P_y}{P_x}" .........(i

"100=P_x(\frac{2P_yY}{P_x})+P_yY"

"100=2P_yY+P_yY"

"100=P_yY(2+1)"


"P_yY=\frac{100}{3}"


"Y^*=\frac{100}{3P_y}...... Demand function for Y"


"100=P_xX+P_y(\frac{0.5P_xX}{P_y})"

"P_xX=\frac{100}{1.5}"


"X^*=\frac{100}{1.5P_x}".......Demand function of X


b) Maximizing Levels of X and Y

"\frac{5x^\frac{-1}{2}Y^\frac{1}{4}}{2.5x^\frac{1}{2}Y^\frac{-3}{4}}=\frac{4}{2}"


"\frac{5Y}{2.5X}=\frac{4}{2}"


Y=X


"P_x\times X= P_y\times Y= m"


"4Y+2Y=100"


"Y=\frac{100}{6}=\frac{50}3"


X= "\frac{50}3"


c) Maximum Utility= "\frac {50}{3}+\frac{50}{3}=\frac{100}{3}=33.33"


d) Marginal rate of substitution of the X and Y

"MRS_{x_y}= \frac{\delta U}{(\delta_x\delta_y)}"


"MRS_{x_y}= \frac{0.125}{5\times 2.5}" =0.01


"MRS_{y_x}= \frac{0.125}{2.5\times 5}" =0.01

There is no difference in the MRS